Gary is 5 years older than harry. Twenty ninw yrs ago gary was twice as old as harry. Hiw old are they now?
Answer by JulietG(1812)

G = H + 5
G-29 = 2(H-29)
G-29 = 2H-58
Substitute the known value of G from the first equation into the second.
(H+5)-29 = 2H-58
H-24 = 2H-58
-24 = H-58
34 = H
If Harry is 35, then Gary is 39.

Check:
Twenty nine yrs ago gary was (39-29=10) twice as old as harry (35-29=5; 5*2=10)